OpenStudy (anonymous):

tan theta =3/4 find sin theta

OpenStudy (abb0t):

tangent = \(\sf \frac{opposite}{adjacent}\) use pythagorean theorem to find sine, which is: \(\sf \frac{opposite}{hypotenuse}\)

OpenStudy (abb0t):

Pythagorean theorem \(\Rightarrow\) \(\sf a^2 +b^2=c^2\) c = \(\sf \sqrt{a^2+b^2}\)